OpenStudy (anonymous):

If z = arctan(y/x), find d^2z/dxdy

terenzreignz (terenzreignz):

\[\huge z \ = \ \tan^{-1}\left(\frac{y}x\right)\] And you're to find \[\huge \frac{\partial^2z}{\partial x \partial y}\] right?

terenzreignz (terenzreignz):

Partial derivatives... first, find its derivative with respect to y (or x, really, but I think y is better) Find its derivative with respect to y whilst pretending for the moment that x is a constant.

OpenStudy (anonymous):

Yes that's what I meant, sorry didn't see the "Equation" button, woops.

terenzreignz (terenzreignz):

So... can you find the derivative with respect to y?

OpenStudy (anonymous):

\[\frac{1}{ 1 + (\frac{ Y }{ X })^{2}}\] Am I on the right track?

terenzreignz (terenzreignz):

Quite so :) Now take the derivative of THIS expression, this time with respect to x. And THIS time, pretend y is a constant.

OpenStudy (amistre64):

so, in this notation its read from right to left, correct? dy then dx

terenzreignz (terenzreignz):

Are you referring to the order in which this is differentiated?

OpenStudy (anonymous):

I'm having a blank moment :/

terenzreignz (terenzreignz):

At differentiating this function: \[\huge \frac{\partial}{\partial x}\frac{1}{1+\left(\frac{y}x\right)^2}\]?

OpenStudy (anonymous):

Yes

terenzreignz (terenzreignz):

Okay, chain rule... first, the derivative of 1/x, we apply it here... \[\huge -\frac{1}{\left(1+\left(\frac{y}x\right)^2\right)^2}\cdot\frac{\partial}{\partial x}\left(1+\left(\frac{y}{x}\right)^2\right)\] Does this make you feel better? :D

OpenStudy (amistre64):

\[D_y[tan^{-1}(ky)]=\frac{k}{1+(ky)^2}\]

OpenStudy (anonymous):

Much better :D and to get 2nd derivative I just derive again

terenzreignz (terenzreignz):

Yes. The word is differentiate, not derive, by the way :D

OpenStudy (anonymous):

Alright woops my bad, terrible text book this

terenzreignz (terenzreignz):

Lol, that's semantics, this is mathematics. Please proceed :)

OpenStudy (amistre64):

your derivative is missing the constant (1/x) ... unless ive overlooked it

terenzreignz (terenzreignz):

I think the y/x does that?

OpenStudy (amistre64):

no. what is the derivative w.r.t.y of say: arctan(6y) ?

OpenStudy (amistre64):

\[D_y[\tan^{-1}(6y)]\ne\frac{1}{1+(6y)^2}\]

terenzreignz (terenzreignz):

Snap... You're RIGHT @MathsNoob URGENT We forgot the 1/x bit when we differentiated with respect to y!

OpenStudy (amistre64):

it happens :)

terenzreignz (terenzreignz):

This is why I hate calculus... actually, I don't have a reason, I just hate it :D

OpenStudy (amistre64):

see if i can write this up as a product rule now .... \[D_x[x^{-1}~(1+kx^{-2})^{-1}]\]

OpenStudy (amistre64):

such that k=y^2

OpenStudy (anonymous):

So... \[z = \tan ^{-1}(\frac{ Y }{ X })\] \[= (\frac{ 1 }{ 1 +(\frac{ Y }{ X })^{2} }). (-\frac{ Y }{ X ^{2} }) + (\frac{ 1 }{ 1 +(\frac{ Y }{ X })^{2} }). (-\frac{ 1 }{ X })\] ???

OpenStudy (amistre64):

it might prove better to simplify Zy first \[\frac{\frac 1x}{1+\frac{y^2}{x^2}}\] \[\frac{x^2}{x^2}\frac{\frac 1x}{1+\frac{y^2}{x^2}}\] \[\frac{x}{x^2+y^2}\]

OpenStudy (amistre64):

\[D[fg]=f'g+fg'\] \[f=x~:~f'=1\] \[g=(x^2+k)^{-1}~:~g'=-(x^2+k)*2x\]

OpenStudy (amistre64):

-(x^2+k)^(-2) that is :)

OpenStudy (amistre64):

\[Z_{yx}=1(x^2+y^2)^{-1}+x(-2x)(x^2+y^2)^{-2}\] \[Z_{yx}=[(x^2+y^2)+x(-2x)](x^2+y^2)^{-2}\] \[Z_{yx}=\frac{x^2+y^2-2x^2}{(x^2+y^2)^2}\]